OpenStudy (anonymous):

Use transformations to solve the equation. e – 14.4 = –20.8

OpenStudy (anonymous):

35.2?

OpenStudy (fibonaccichick666):

Might I enquire what class this is for? I have not seen that word(transformations) used in this situation before and would like to look it up.

OpenStudy (anonymous):

Its for math class

OpenStudy (anonymous):

e – 14.4 = –20.8; e = -20.8 + 14.4 e = -6.4 i got this

OpenStudy (anonymous):

Thx

OpenStudy (anonymous):

you welcome

OpenStudy (fibonaccichick666):

also, if they just mean solve for e, think about it as performing the order of operations backwards. Also remember addition of negatives and positives

OpenStudy (anonymous):

Yay i got 100 percent

OpenStudy (anonymous):

guuudd